whats the anwser for 5 under2t-t=3+3 under 2t
??????? i need help with the steps

Answer by mananth(16946)

5 under2t-t=3+3 under 2t
5/2t -t = 3 + 3/2t
5/2t+3/2t -t=3
LCM of left side = 2t
(5+3-2t^2) / 2t =3
(5+3-2t^2) =6t
8-2t^2=6t
2t^2+6t-8=0 ( divide by 2)
t^2+3t-4=0
t^2-4t+t-4=0
t(t-4)+1(t-4)=0
(t-4)(t+1)=0
t=4 OR -1